Activity One: Given A Joint Frequency Table Location of adverstised position Outside Sydney Sydney-based Total Sector Management Marketing Accounting Finance Economics Other 19 14 14 15 18 31 30 46 29 42 41 81 269 49 60 43 57 59 112 380 Total a) What is the probability the position will be in Finance? b) What is the probability the position will be Sydney-based? c) What is the probability the position will be in Finance and located in Sydney? d) What is probability an advertised position will be in Finance, given that it is a Sydney-based position? What is probability an advertised position will be Sydney-based, given that th position is in Finance? e)
Answer #1

a)

Out of 380 positions, 57 are in finance so

P(finance) = 57 /380 = 0.15

b)

Out of 380 positions, 269 are Sydney based so

P(Sydney based) = 269 /380 = 0.7079

c)

P(Finance and Sydney based) = 42 /380 = 0.1105

d)

P(Finance | Sydney based) = P(Finance and Sydney based) / P(Sydney based) = 42 / 269 = 0.1561

e)

P(Sydney based | Finance) = P(Finance and Sydney based) / P(finance) = 42 / 57 = 0.7368